Operation Summary for March 4, 2011 - March 7, 2011 |
The Integrated Luminosity for the period from 2/28/11 to 3/7/11 was 52.1 pb-1 (inverse picobarns). NuMI was off to replace its target. |
|
Friday March 4 The day shift began with Operations monitoring the Tevatron (TEV) store 8547 (100.63E30) and the Recycler stash (290.1E10), with the Antiproton Source (Pbar) stacking (9.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and FTBF T-1012 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8547 at 12:40 PM, put the TEV into HEP shot setup, and established store 8548 at 1:53 PM with an average initial luminosity (AIL) of 347.47E30. Operations reported some downtime when a Linac RF (LRF5) console was rebooted. This reboot took out the Linac low-level RF for about 30 minutes. During the evening shift, Operations disabled beam to T-1012 at 6:15 PM. Operations reported a power outage in the G2 service building (a non-operating Switchyard beam line) due to a tripped breaker in the building; the Duty Electrician reset the breaker.Saturday March 5 The midnight shift began with Operations monitoring the store 8548 (103.4E30) and the stash (269.0E10), with Pbar stacking (15.1E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8548 at 4:53 AM, put the TeV into shot setup, and established store 8549 at 6:07 AM with an AIL of 333E30. Operations established beam to T-1012 at 6:43 AM. During the day shift, Operators monitored the accelerator complex. During the evening shift, operators disabled beam to T-1012 at 6 PM. Operations terminated store 8549 at 7:03 PM and put the TEV into shot setup, but the TEV quenched when beam was injected on top of beam already in the machine due to the abort kickers not charging and firing. Recovery followed. Operators reduced the beam intensity through Booster in anticipation of tunnel work during the shutdown. The Switchyard F Septa tripped off due to vacuum problems, but since T-1012 had already turned off, this wasn’t a problem. Vacuum personnel will look at the septa vacuum during the shutdown. Operations reported that the TEV sector E48 injection kicker tripped off on a low insulating gas (SF6) indication; operators filled the system with gas and got the kicker to reset. Operations reported that the TEV had recovered and operators put the TEV through a dry squeeze and a wet squeeze. Operations put the TEV into shot setup at 10:38 PM, but the TEV abort kickers again failed to charge up, and while trying to recover from this the injection kicker again fired into the existing beam, which caused another quench. Operators reported that a TEV high-order corrector magnet power supply (T:OD) would not reset after the quench. Operators called in EE Support techs to investigate.Sunday March 6 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations reported that the recovery from this quench was very quick since there was very little beam in the machine. Operators consulted with kicker specialists, and changed out a CAMAC 377 kicker timer card. Operators called in RF technicians to repair a TEV RF station (TRF3) that developed a LCW flow sensor problem after the quench. Operations reported some downtime due to LRF3 trips. Operators had the TEV in a dry squeeze at the end of the shift. During the day shift, Operations put the TEV into shot setup at 9:14 AM and established store 8556 at 10:16 AM with an AIL of 368.68E30. During the evening shift, Operations terminated store 8556 at 6:51 PM and put the TEV into shot setup, but the TEV quenched near the end of scraping. A MiniBooNE LCW skid shut itself down twice; operators restarted the skid each time. Operations reported the output of the I- Source dropped off early in the shift; operators and Linac personnel were able to tune the source and recover the lost output. Operations halted stacking at 7:21 PM for studies. Operators searched and secured the following areas: G2, N01, NM2, and New Muon Lab. Operators reported that they could find no cause for the TEV quench, other than the horizontal antiproton and proton emittances had blown up at the start of scraping. The TEV cryo system recovery was completed at 11:43 PM.Monday March 7 The midnight shift began with the TEV in a dry squeeze, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, TEV personnel conducted studies. Pbar personnel conducted studies. Operators prepared Pbar for access. Recycler personnel conducted studies. Operations turned off beam to the TEV, MI, MiniBooNE, and Booster in preparation for the upcoming shutdown. The Plans for Monday The plans for today are to conduct upgrades and necessary repairs to the accelerator complex while ComEd conducts mandated high-voltage tests. |
|
More Information Comments and Suggestions |
| last modified 03/07/2011 email Fermilab |
| Security, Privacy, Legal | |